# New Insights Into Achievable Spectral Efficiency With Adaptive Free Space Optical Transmissions

### v8

@@ II-B Noise Sources in FSO Wireless Systems
The IM-DD photodetector output consists of the desirable photocurrent component mapped one-to-one to the intensity of the incoming (information) optical signal, signal-dependent shot noise, ambient shot noise, and thermal noise currents. The operating mode of the IM-DD receiver is determined by the relative strengths of these noise components, as follows.
  • For an IM-DD receiver operating in intense ambient light conditions and/or strong thermal noise conditions, it is reasonable that the corrupting noise is signal-independent. It typically holds for indoor IM-DD infrared diffuse or non-LOS links [43].
  • For the IM-DD based FSO systems with large received optical information-signal powers, it is reasonable to expect the corrupting noise to be signal-dependent [44]. It is valid for short-haul and/or medium-haul laser-based IM-$D$ LOS directed links with reasonable transmit powers [43].
In a coherent HD receiver, an optical local oscillator (OLO) signal is mixed with the received signal prior to photodetection. This introduces an OLO-dependent shot noise current in addition to signal-dependent shot noise, ambient shot noise, and thermal noise currents. By ensuring a large OLO optical power, the OLO shot noise component can be made dominant compared to other noise components; this is in fact considered to be the de facto mode of operation of an (ideal) coherent HD receiver [10][42][45].
@@
II-C Receiver Models
Let $x_{\mathrm{\scriptscriptstyle IMDD}}\in\mathbb{R}^{+}$ and $x_{\mathrm{\scriptscriptstyle HD}}\in\mathbb{C}$ denote the transmitted signals for the IM-DD and coherent HD systems respectively. We remind the reader that the baseband (information) signal in a coherent HD scheme represents a complex optical field, whereas the baseband signal in an IM-DD scheme represents optical intensity. In both schemes, the transmitted beam power is constrained by a fixed average power constraint denoted by: $\mathbb{E}[x_{\mathrm{\scriptscriptstyle IMDD}}]\leqslant P_{\mathrm{avg}}$ and $\mathbb{E}\left[|x_{\mathrm{\scriptscriptstyle HD}}|^{2}\right]\leqslant P_{\mathrm{avg}}$. As justified in the introduction, a frequency-flat block fading model is well-suited to capture the random irradiance fluctuations in a terrestrial optical channel: each fading state is assumed to span an identical duration, corresponding to a fixed coherence time, with i.i.d. fading realizations across blocks. The instantaneous channel state is considered to be perfectly known at both the transmitter and the receiver at all times. In practice, this state needs to be estimated only once at the beginning of each block. Within this framework, we consider a FSO transmitter that performs signal adaptation on a block-by-block basis. Adhering to this understanding, the received signal models for the two detection systems are developed next.
1) Signal-Dependent IM-DD Receiver System: The baseband signal in an IM-DD receiver is described by
$y_{\mathrm{\scriptscriptstyle IMDD}}\,=\,Ix_{\mathrm{\scriptscriptstyle IMDD}}\,\,+\,\,\sqrt{Ix_{\mathrm{\scriptscriptstyle IMDD}}}\,\,n_{\mathrm{\scriptscriptstyle SD}}\,\,+\,\,n_{\mathrm{\scriptscriptstyle SI}}$ (8)
where $y_{\mathrm{\scriptscriptstyle IMDD}}$ is the received signal, $I$ is the intensity fluctuation state of the terrestrial FSO link (see (7)), $n_{\mathrm{\scriptscriptstyle SI}}\sim\mathcal{N}(0,\sigma_{\mathrm{\scriptscriptstyle SI}}^{2})$ is the signal-independent noise (combined ‘ambient and thermal noise’) modeled as additive white Gaussian noise (AWGN), and the $\sqrt{Ix_{\mathrm{\scriptscriptstyle IMDD}}}\,n_{\mathrm{\scriptscriptstyle SD}}\sim\mathcal{N}(0,Ix_{\mathrm{\scriptscriptstyle IMDD}}\,\sigma^{2})$ is the signal-dependent AWGN [44].
Remark 2.
In the signal-independent noise-limited IM-DD operation, $n_{\mathrm{\scriptscriptstyle SI}}>>\sqrt{Ix_{\mathrm{\scriptscriptstyle IMDD}}}n_{\mathrm{\scriptscriptstyle SD}}$ with high probability. Hence, the received SNR $(:=I^{2}x^{2}_{\mathrm{\scriptscriptstyle IMDD}}/\sigma_{\mathrm{\scriptscriptstyle SI}}^{2})$ varies as the ‘square’ of the received optical power. In [25], [26], [27], the average SE maximization problem for the signal-independent noise-limited IM-DD channel is formulated as $\mathbb{E}[\ln(1+I^{2}P^{2}(I))]$ where $P^{2}(I):=x^{2}_{\mathrm{\scriptscriptstyle IMDD}}$ under the constraint $\mathbb{E}[P(I)]=P_{\mathrm{avg}}$; this is a non-convex optimization, and more seriously, the above SE formulation ignored the non-negativity constraint on $I$ (see [48] for details).
In this work, we consider the signal-dependent noise-limited IM-DD receiver; thus, the noise power is dominated by $Ix_{\mathrm{\scriptscriptstyle IMDD}}\,\sigma^{2}$. Hence, conditioned on $I$, the received SNR is
$\gamma_{\mathrm{\scriptscriptstyle IMDD}}\,=\,{Ix_{\mathrm{\scriptscriptstyle IMDD}}}/{\sigma^{2}}.$ (9)
Notice from (9) that the SNR varies linearly with the incident optical power at the considered IM-DD receiver.
2) Coherent Heterodyne Detection Receiver System: The complex baseband signal in a coherent HD receiver is given by
@@ II-C Receiver Models
where $y_{\mathrm{\scriptscriptstyle HD}}\ \in\ \mathbb{C}$ represents its output, $h\in\mathbb{C}$ the fluctuation in the received optical field, $w_{\mathrm{\scriptscriptstyle OLO}}\sim\mathcal{CN}(0,\sigma^{2}_{\mathrm{\scriptscriptstyle OLO}})$ the OLO shot noise, and $w_{\mathrm{\scriptscriptstyle OTH}}$ the total AWGN from all the sources other than the OLO [13, Ch. 3]. The OLO shot noise is the only dominant noise of consequence. Hence, conditioned on $h$ or more appropriately on $I:=|h|^{2}$, the received SNR is
@@ III Achievable Spectral Efficiency With Adaptive Optical Beam Transmission
The transmitted signal $x_{\mathrm{\scriptscriptstyle IMDD}}$, representing optical intensity, must be non-negative. In contrast, the $x_{\mathrm{\scriptscriptstyle HD}}$ scheme has no such constraint and can utilize two degrees of freedom. As a result, the classical Shannon limit on spectral efficiency for the AWGN channel applies to the HD scheme, but not to the IM-DD system (see [48]). Conditioned on the received SNRs in (9) and (11), the SE of the optical channels (expressed in nats/s/Hz) based on IM-DD and HD schemes satisfy
$S_{\mathrm{\scriptscriptstyle IMDD}}(I)\,$ $\geq\,\dfrac{1}{2}\ln\left(1+\dfrac{\gamma_{E}}{2\pi}\dfrac{Ix_{\mathrm{\scriptscriptstyle IMDD}}}{\sigma^{2}}\right),$ (12)
$S_{\scriptscriptstyle\mathrm{HD}}(I)\,$ $=\,\,\ln\left(1+\dfrac{I\,|x_{\mathrm{\scriptscriptstyle HD}}|^{2}}{\sigma_{\mathrm{\scriptscriptstyle OLO}}^{2}}\right).$ (13)
In (12), the constant $\gamma_{E}$ is the Euler’s constant. The lower bound on the achievable SE of the signal-dependent IM-DD channel, given by (12), is derived from [48, Theorem 7] for the AWGN IM-DD channel under an average-power constraint. This bound approaches the actual SE at high received optical powers [48, Proposition 8], making it a reliable approximation in such regimes. Using (12) and (13), a unified SE expression for a given channel state $\lambda:=I$ can be derived for both FSO systems as follows.
$S(\lambda)\,:=\,\frac{1}{k}\ln\left(1+\frac{\lambda P_{k}(\lambda)}{N_{k}}\right)$ (14)
where $k=1$ for the HD based scheme with $N_{1}={\sigma^{2}_{\mathrm{\scriptscriptstyle OLO}}}\,,\,P_{1}(\lambda):=|x_{\mathrm{\scriptscriptstyle HD}}|^{2}$, and $k=2$ for the IM-DD based scheme with $N_{2}=\sigma^{2}/\left({\gamma_{E}}/{2\pi}\right),\,P_{2}(\lambda):=x_{\mathrm{\scriptscriptstyle IMDD}}$. The power constraint mentioned earlier for both schemes gets translated to $\mathbb{E}[P_{k}(\lambda)]\leqslant P_{\mathrm{avg}}$.
The average (long-term) spectral efficiency (ASE) in nats/s/Hz is given by
$\frac{1}{k}\,\mathbb{E}\left[\ln\left(1+\frac{\lambda P_{k}(\lambda)}{N_{k}}\right)\right]$ (15)
where the maximization is performed over all feasible transmit laser beam adaptation schemes that satisfy $\mathbb{E}[P_{k}(\lambda)]\leqslant P_{\mathrm{avg}}$.
Note that ${P_{\mathrm{avg}}}/{N_{k}}$ represents the average transmit power normalized by the average noise power, which will henceforth be termed the transmit SNR and denoted by $\mathrm{SNR}_{k}$. The optimal beam power adaptation $P_{k}^{\,*}$ maximizing (LABEL:eq:capacity) is
$\dfrac{P_{k}^{\,*}(\lambda)}{N_{k}}\,=\,\left(\dfrac{1}{\mu_{k}}-\dfrac{1}{\lambda}\right)^{+}$ (16)
with $\mu_{k}$ chosen so that the optical power constraint
$\mathbb{E}\biggl{[}\left(\dfrac{1}{\mu_{k}}-\dfrac{1}{\lambda}\right)^{+}\biggr{]}\,=\,\mathrm{SNR}_{k}$ (17)
is satisfied. Additionally, the parameter $\mu_{k}$ serves as a channel cutoff, which simplifies the ASE problem in (LABEL:eq:capacity) to
@@
$\phantom{xxxxx}\underbrace{\int_{0}^{\mu_{k}}\ln(\lambda)f_{\lambda}(\lambda)d\lambda\ }_{=\,I_{2}}\,-\,\underbrace{\int_{\mu_{k}}^{\infty}\ln(\mu_{k})f_{\lambda}(\lambda)d\lambda}_{=\,I_{3}}\,\,\Biggr{]}.$ (19)
@@
$G^{m,\,n+1}_{p+1,\,q+1}\hskip-2.5pt\left(x\,\middle|\hskip-3.50006pt\begin{array}[]{c}1,j\!+a_{1},\ldots,j\!+a_{n},j\!+a_{n+1},\ldots,j\!+a_{p}\\ j\!+b_{1},\ldots,j\!+b_{m},0,j\!+b_{m+1},\ldots,j\!+b_{q}\end{array}\hskip-6.49994pt\right).$ (25)
@@ III-A Exact ASE With and Without Pointing Errors
For the GG turbulence channel with pointing error and perfect channel state information (CSI) at both ends, the exact ASE is given by
@@
$\phantom{xxxxxx}\,\,+\,\frac{\xi^{2}}{\Gamma(a)\Gamma(b)}G_{3,5}^{3,2}\left(\frac{ab}{A_{0}}\mu_{k}\,\biggr{\rvert}\hskip-1.49994pt\begin{array}[]{c}1,\,1,\,\xi^{2}+1\\[1.00006pt] \xi^{2},\,a,\,b,\,0,\,0\\ \end{array}\hskip-5.0pt\right)\,\biggr{]}$ (28)
where $\psi(\cdot)$ is the Euler’s digamma function, and $\mu_{k}$ is determined from $\int_{\mu_{k}}^{\infty}(1/\mu_{k}-1/\lambda)f_{{\lambda}}(\lambda)d\lambda\,=\,\mathrm{SNR}_{k}$.
@@
The unified ASE expression (from (III)) is
@@
$\frac{ab\xi^{2}}{A_{0}\Gamma(a)\Gamma(b)}\int_{0}^{\infty}\lambda^{t}\,G_{1,3}^{3,0}\left(\frac{ab\lambda}{A_{0}}\biggr{\rvert}\hskip-2.5pt\begin{array}[]{c}\xi^{2}\\ \xi^{2}-1,a-1,b-1\\ \end{array}\hskip-5.0pt\right)d\lambda.$ (34)
Applying (22) in (III-A) along with the substitutions that
$\lim_{x\to 0}\,\,\,$ $G_{2,4}^{3,1}\left(x\,\biggr{\rvert}\hskip-1.49994pt\begin{array}[]{c}1,\,\xi^{2}+t+1\\ \xi^{2}+t,\,a+t,\,b+t,\,0\\ \end{array}\hskip-5.0pt\right)\,=\,0,\,\,\,\text{and}$
$\lim_{x\to\infty}\,\,$ $G_{2,4}^{3,1}\left(x\,\biggr{\rvert}\hskip-1.49994pt\begin{array}[]{c}1,\,\xi^{2}+t+1\\ \xi^{2}+t,\,a+t,\,b+t,\,0\\ \end{array}\hskip-5.0pt\right)\,=\,\frac{\Gamma\left(a+t\right)\Gamma\left(b+t\right)}{\xi^{2}+t},$
we have
$I_{1}\,=\,\frac{\xi^{2}}{\Gamma(a)\Gamma(b)}\frac{d}{dt}\left(\left(\frac{A_{0}}{ab}\right)^{t}\frac{\Gamma\left(a+t\right)\Gamma\left(b+t\right)}{\xi^{2}+t}\right)\Biggl{|}_{t=0}.$ (35)
@@
$I_{1}\ =\ \ln\left(\frac{A_{0}}{ab}\right)+\psi(a)+\psi(b)-\left(\frac{1}{\xi^{2}}\right)\,\cdot$ (36)
@@
$I_{2}=\int_{0}^{\mu_{k}}\ln(\lambda)f_{\lambda}(\lambda)d\lambda\phantom{xxxxxxxxxxxxxxxssssssssssssxx}\,$
$=\frac{ab\xi^{2}}{A_{0}\Gamma(a)\Gamma(b)}\int_{0}^{\mu_{k}}\ln\left(\lambda\right)G_{1,3}^{3,0}\left(\hskip-1.00006pt\frac{ab\lambda}{A_{0}}\biggr{\rvert}\hskip-4.49997pt\begin{array}[]{c}\xi^{2}\\[1.00006pt] \xi^{2}-1,a-1,b-1\\ \end{array}\hskip-5.50003pt\right)\hskip-1.49994ptd\lambda.$ (39)
@@
$I_{2}\,$ $=\,\biggl{[}\,G_{2,4}^{3,1}\left(\dfrac{ab\lambda}{A_{0}}\,\biggr{\rvert}\hskip-1.49994pt\begin{array}[]{c}1,\xi^{2}+1\\[1.00006pt] \xi^{2},a,b,0\end{array}\hskip-5.0pt\right)\times\ln(\lambda)$ (42)
$\,\,\,\,\,\,\,\,\,\,\,\,\,\,-G_{3,5}^{3,2}\left(\dfrac{ab\lambda}{A_{0}}\,\biggr{\rvert}\hskip-1.49994pt\begin{array}[]{c}1,1,\xi^{2}+1\\[1.00006pt] \xi^{2},a,b,0,0\end{array}\hskip-5.0pt\right)\,\biggr{]}\,\biggl{|}_{\lambda=0}^{\lambda=\mu_{k}}\,\,\times\dfrac{\xi^{2}}{\Gamma(a)\Gamma(b)}.$ (45)
With the Meijer-G function limits
$\lim_{x\to 0}\,\,$ $G_{2,4}^{3,1}\left(x\,\biggr{\rvert}\hskip-1.49994pt\begin{array}[]{c}1,\xi^{2}+1\\[1.00006pt] \xi^{2},a,b,0\end{array}\hskip-5.0pt\right)\,\times\,\ln(x)\,=\,0,\,\,\,\text{and}$
$\lim_{x\to 0}\,\,$ $G_{3,5}^{3,2}\left(x\,\biggr{\rvert}\hskip-1.49994pt\begin{array}[]{c}1,1,\xi^{2}+1\\[1.00006pt] \xi^{2},a,b,0,0\end{array}\hskip-5.0pt\right)\,=\,0,$
@@
$I_{2}\,$ $=\,\left[\,G_{2,4}^{3,1}\left(\dfrac{ab}{A_{0}}\mu_{k}\,\biggr{\rvert}\hskip-1.49994pt\begin{array}[]{c}1,\xi^{2}+1\\[1.00006pt] \xi^{2},a,b,0\end{array}\hskip-5.0pt\right)\right.\times\ln(\mu_{k})$ (48)
$\,\,\,-\left.G_{3,5}^{3,2}\left(\dfrac{ab}{A_{0}}\mu_{k}\,\biggr{\rvert}\hskip-1.49994pt\begin{array}[]{c}1,1,\xi^{2}+1\\[1.00006pt] \xi^{2},a,b,0,0\end{array}\hskip-5.0pt\right)\,\right]\dfrac{\xi^{2}}{\Gamma(a)\Gamma(b)}\cdot$ (51)
@@
$I_{3}\,$ $=\,\ln(\mu_{k})\int_{\mu_{k}}^{\infty}f_{\lambda}(\lambda)d\lambda,$ (52)
@@
$\int_{\mu_{k}}^{\infty}f_{\lambda}(\lambda)d\lambda\phantom{xxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxx}$
= $=\,1-\frac{ab\xi^{2}}{A_{0}\Gamma(a)\Gamma(b)}\int^{\mu_{k}}_{0}G_{1,3}^{3,0}\left(\frac{ab\lambda}{A_{0}}\,\biggr{\rvert}\hskip-1.49994pt\begin{array}[]{c}\xi^{2}\\[1.00006pt] \xi^{2}-1,a-1,b-1\\ \end{array}\hskip-5.0pt\right)d\lambda$ (55)
= $=\,1-\frac{\xi^{2}}{\Gamma(a)\Gamma(b)}G_{2,4}^{3,1}\left(\frac{ab}{A_{0}}\mu_{k}\,\biggr{\rvert}\hskip-1.49994pt\begin{array}[]{c}1,\,\xi^{2}+1\\[1.00006pt] \xi^{2},\,a,\,b,\,0\\ \end{array}\hskip-5.0pt\right).$ (58)
@@
$\lim_{x\to 0}\,\,G_{2,4}^{3,1}\left(x\,\biggr{\rvert}\hskip-1.49994pt\begin{array}[]{c}1,\,\xi^{2}+1\\[1.49994pt] \xi^{2},\,a,\,b,\,0\\ \end{array}\hskip-5.0pt\right)\,=\,0.$
@@
For the GG turbulence channel ‘without pointing error’ and perfect CSI at both ends, the exact ASE is given by
@@
$\phantom{xxxxxx}+\,\frac{1}{\Gamma(a)\Gamma(b)}G_{2,4}^{2,2}\left({ab}\mu_{k}\,\biggr{\rvert}\hskip-1.49994pt\begin{array}[]{c}1,\,1\\[1.00006pt] a,\,b,\,0,\,0\\ \end{array}\hskip-5.0pt\right)\,\biggr{]}$ (61)
where $\mu_{k}$ is solved from $\int_{\mu_{k}}^{\infty}(1/\mu_{k}-1/\lambda)f_{{\lambda}}(\lambda)d\lambda\,=\,\mathrm{SNR}_{k}$.

### v9

@@ II-B Noise Sources
The IM-DD photodetector output consists of the desirable photocurrent component mapped one-to-one to the intensity of the incoming (information) optical signal, signal-dependent shot noise, ambient shot noise, and thermal noise currents. The operating mode of an IM-DD receiver is determined by the relative strengths of these noise components.
In a coherent HD receiver, in contrast, an optical local oscillator (OLO) signal is mixed with the received signal prior to photodetection. This introduces an OLO-dependent shot noise current in addition to signal-dependent shot noise, ambient shot noise, and thermal noise currents. By ensuring a large OLO optical power, the OLO shot noise component can be made dominant compared to other noise components; this is in fact considered to be the de facto mode of operation of an (ideal) coherent HD receiver [10][42][43].
@@
II-C Coherent HD Receiver Model
Let $x_{\mathrm{\scriptscriptstyle HD}}\in\mathbb{C}$ denote the transmitted symbol for the coherent HD system respectively. We remind the reader that the baseband (information) signal in a coherent HD scheme represents a complex optical field, whereas the baseband signal in an IM-DD scheme represents optical intensity. The transmitted beam power is constrained by a fixed average power constraint denoted by $\mathbb{E}\left[|x_{\mathrm{\scriptscriptstyle HD}}|^{2}\right]\leqslant P_{\mathrm{avg}}$. As justified in the introduction, owing to millisecond-scale coherence times and hundreds of gigahertz coherence bandwidths, a frequency-flat block fading model is well-justified for a terrestrial optical fading channel subjected to time-varying irradiance fluctuations: each fading state is assumed to span an identical duration, corresponding to a fixed coherence time, with i.i.d. fading realizations across blocks. The block-wise channel state is considered to be perfectly known at both the transmitter and the receiver at all times. In practice, this state needs to be estimated at the receiver only once at the beginning of each block, and conveyed to the transmitter over an error-free feedback link. Within this framework, we consider a FSO transmitter that performs signal adaptation on a block-by-block basis. Adhering to this understanding, the received signal model for the HD detection-based FSO receiver system is developed next.
The complex baseband symbol in a coherent HD receiver is described by
@@ II-C Coherent HD Receiver Model
where $y_{\mathrm{\scriptscriptstyle HD}}\ \in\ \mathbb{C}$ represents its output, $h\in\mathbb{C}$ the fluctuation in the received optical field, $w_{\mathrm{\scriptscriptstyle OLO}}\sim\mathcal{CN}(0,\sigma^{2}_{\mathrm{\scriptscriptstyle OLO}})$ the OLO shot noise, and $w_{\mathrm{\scriptscriptstyle OTH}}$ the total AWGN from all the sources other than the OLO [13, Ch. 3]. As noted earlier, the OLO shot noise is the only dominant noise of consequence. Hence, conditioned on $h$ or more appropriately on $I:=|h|^{2}$, the received SNR is
@@ III Achievable Spectral Efficiency With Adaptive Optical Beam Transmission
In an IM-DD FSO system, we notice that the transmitted signal represents optical intensity, and thus must be non-negative. On the other hand, the $x_{\mathrm{\scriptscriptstyle HD}}$ scheme has no such constraint and can utilize two degrees of freedom. As a result, the classical Shannon limit on spectral efficiency for the AWGN channel applies to the HD scheme, but not to the IM-DD system (see [46]). Conditioned on the received SNR in (11), the achievable SE of the coherent optical channel (in nats/s/Hz) based on HD scheme satisfy
$\mathrm{S}_{\scriptscriptstyle\mathrm{HD}}(\lambda)\,$ $=\,\,\ln\left(1+\dfrac{\lambda P(\lambda)}{N_{\mathrm{hd}}}\right)$ (12)
where $\lambda:=I$, $P(\lambda):=|x_{\mathrm{\scriptscriptstyle HD}}|^{2}$ and $N_{\mathrm{hd}}:=\sigma_{\mathrm{\scriptscriptstyle OLO}}^{2}$. The achievable long-term (over blocks) or average spectral efficiency (ASE) is given by
[…]
where the maximization is performed over all feasible beam adaptations choices of $P$ that satisfy $\mathbb{E}[P(\lambda)]\leqslant P_{\mathrm{avg}}$.
Note that ${P_{\mathrm{avg}}}/{N_{\mathrm{hd}}}$ represents the average transmit power normalized by the average noise power, which will henceforth be termed the transmit SNR and denoted by $\mathrm{SNR}$. The optimal beam power adaptation $P^{\,*}$ maximizing (13) is
$\dfrac{P^{\,*}(\lambda)}{N_{\mathrm{hd}}}\,=\,\left(\dfrac{1}{\mu}-\dfrac{1}{\lambda}\right)^{+}$ (14)
with $\mu$ chosen so that the optical power constraint
$\mathbb{E}\biggl[\left(\dfrac{1}{\mu}-\dfrac{1}{\lambda}\right)^{+}\biggr]\,=\,\mathrm{SNR}$ (15)
is satisfied. Additionally, the parameter $\mu$ serves as a channel cutoff, which simplifies the ASE problem in (13) to
@@
$\phantom{xxxxx}\underbrace{\int_{0}^{\mu}\!\!\ln(\lambda)f_{\lambda}(\lambda)d\lambda\ }_{=\,I_{2}}\,-\,\underbrace{\int_{\mu}^{\infty}\!\!\ln(\mu)f_{\lambda}(\lambda)d\lambda}_{=\,I_{3}}\,\,\Biggr].$ (17)
@@
$G^{m,\,n+1}_{\!p+1,\,q+1}\hskip-2.5pt\left(x\middle|\hskip-3.50006pt\begin{array}[]{c}1,j\!+a_{1},\ldots,j\!+a_{n},j\!+a_{n+1},\ldots,j\!+a_{p}\\ j\!+b_{1},\ldots,j\!+b_{m},\!0,j\!+b_{m+1},\ldots,j\!+b_{q}\end{array}\hskip-5.0pt\right)\!.$ (23)
@@ III-A Exact ASE With and Without Pointing Errors
For the coherent HD optical channel subjected to GG turbulence with pointing errors as described in (9), and with perfect channel state information (CSI) at both ends, the exact ASE in (13) is simplified to
@@
$\phantom{x}\,\,\,+\,\frac{\xi^{2}}{\Gamma(a)\Gamma(b)}\,G_{3,5}^{3,2}\left(\frac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\,1,\,\xi^{2}+1\\[1.00006pt] \xi^{2},\,a,\,b,\,0,\,0\\ \end{array}\hskip-5.0pt\right)\biggr]$ (26)
where $\psi(\cdot)$ is the Euler’s digamma function, and $\mu$ is determined from $\int_{\mu}^{\infty}(1/\mu-1/\lambda)f_{{\lambda}}(\lambda)d\lambda\,=\,\mathrm{SNR}$.
@@
The exact ASE expression (from (III)) is
@@
$\frac{ab\xi^{2}}{A_{\mathrm{\scriptscriptstyle 0}}\Gamma(a)\Gamma(b)}\int_{0}^{\infty}\!\!\lambda^{t}G_{1,3}^{3,0}\left(\frac{ab\lambda}{A_{\mathrm{\scriptscriptstyle 0}}}\biggr\rvert\hskip-2.5pt\begin{array}[]{c}\xi^{2}\\ \xi^{2}\!-1,\!a-1,\!b-1\end{array}\hskip-4.49997pt\right)d\lambda.$ (32)
Applying (20) in (III-A), and making the substitutions:
$\phantom{xx}\lim_{x\to 0}\,G_{2,4}^{3,1}\!\left(x\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\xi^{2}\!+t+1\\ \!\xi^{2}\!+t,\!a\!+t,\!b\!+t,\!0\\ \end{array}\hskip-5.0pt\right)=0,\,\,\,\text{and}$ (35)
$\phantom{xx}\lim_{x\to\infty}\,G_{2,4}^{3,1}\!\left(x\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\xi^{2}\!+t+1\\ \!\xi^{2}\!+t,\!a\!+t,\!b\!+t,\!0\\ \end{array}\hskip-5.0pt\right)=\frac{\Gamma\!\left(a\!+t\right)\!\Gamma\!\left(b\!+t\right)}{\xi^{2}\!+t},$ (38)
we get
$I_{1}\,=\,\frac{\xi^{2}}{\Gamma\!(a)\Gamma\!(b)}\,\frac{d}{dt}\left(\left(\frac{A_{\mathrm{\scriptscriptstyle 0}}}{ab}\right)^{t}\frac{\Gamma\left(a+t\right)\Gamma\left(b+t\right)}{\xi^{2}+t}\right)\Biggl|_{\,t=0}.$ (39)
@@
$I_{1}\ =\ \ln\left(\frac{A_{\mathrm{\scriptscriptstyle 0}}}{ab}\right)+\psi(a)+\psi(b)-\left(\frac{1}{\xi^{2}}\right)\,\cdot$ (40)
@@
$I_{2}=\int_{0}^{\mu}\!\!\ln(\lambda)f_{\lambda}(\lambda)d\lambda\phantom{xxxxxxxxxxxxxxxssssssssssssxx}\,$
$=\frac{ab\xi^{2}}{A_{\mathrm{\scriptscriptstyle 0}}\Gamma(a)\Gamma(b)}\int_{0}^{\mu}\ln\left(\lambda\right)G_{1,3}^{3,0}\left(\hskip-1.00006pt\frac{ab\lambda}{A_{\mathrm{\scriptscriptstyle 0}}}\biggr\rvert\hskip-4.49997pt\begin{array}[]{c}\xi^{2}\\[1.00006pt] \xi^{2}-1,a-1,b-1\\ \end{array}\hskip-5.50003pt\right)\hskip-1.49994ptd\lambda.$ (43)
@@
$I_{2}\,$ $=\,\biggl[G_{2,4}^{3,1}\left(\dfrac{ab\lambda}{A_{\mathrm{\scriptscriptstyle 0}}}\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\xi^{2}+1\\[1.00006pt] \xi^{2},a,b,0\end{array}\hskip-3.50006pt\right)\times\ln(\lambda)$ (46)
$\,\,\,\,\,\,\,\,\,-G_{3,5}^{3,2}\left(\dfrac{ab\lambda}{A_{\mathrm{\scriptscriptstyle 0}}}\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,1,\xi^{2}+1\\[1.00006pt] \xi^{2},a,b,0,0\end{array}\hskip-3.50006pt\right)\biggr]\,\biggl|_{\,\lambda=0}^{\,\lambda=\mu}\,\,\times\,\,\,\dfrac{\xi^{2}}{\Gamma(a)\Gamma(b)}.$ (49)
With the Meijer-G function limits:
$\phantom{xxxxxx}\lim_{x\to 0}\,\,G_{2,4}^{3,1}\left(x\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\xi^{2}+1\\[1.00006pt] \xi^{2},a,b,0\end{array}\hskip-3.50006pt\right)\,\times\,\ln(x)\,=\,0,\,\,\,\text{and}$ (52)
$\phantom{xxxxxx}\lim_{x\to 0}\,\,G_{3,5}^{3,2}\left(x\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,1,\xi^{2}+1\\[1.00006pt] \xi^{2},a,b,0,0\end{array}\hskip-3.50006pt\right)\,=\,0,$ (55)
@@
$I_{2}\,$ $=\,\left[G_{2,4}^{3,1}\left(\dfrac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\xi^{2}+1\\[1.00006pt] \xi^{2},a,b,0\end{array}\hskip-3.50006pt\right)\right.\times\,\ln(\mu)$ (58)
$\,\,\,\,\,\,\,\,\,-\left.G_{3,5}^{3,2}\left(\dfrac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,1,\xi^{2}+1\\[1.00006pt] \xi^{2},a,b,0,0\end{array}\hskip-3.50006pt\right)\right]\dfrac{\xi^{2}}{\Gamma(a)\Gamma(b)}\cdot$ (61)
@@
$I_{3}\,$ $=\,\ln(\mu)\int_{\mu}^{\infty}\!\!f_{\lambda}(\lambda)d\lambda,$ (62)
@@
$\int_{\mu}^{\infty}\!\!f_{\lambda}(\lambda)d\lambda\phantom{xxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxx}$
= $=\,1-\frac{ab\xi^{2}}{A_{\mathrm{\scriptscriptstyle 0}}\Gamma(a)\Gamma(b)}\int^{\mu}_{0}\!\!G_{1,3}^{3,0}\left(\frac{ab\lambda}{A_{\mathrm{\scriptscriptstyle 0}}}\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}\xi^{2}\\[1.00006pt] \xi^{2}-1,a-1,b-1\\ \end{array}\hskip-5.0pt\right)d\lambda$ (65)
= $=\,1-\frac{\xi^{2}}{\Gamma(a)\Gamma(b)}\,G_{2,4}^{3,1}\left(\frac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\,\xi^{2}+1\\[1.00006pt] \xi^{2},\,a,\,b,\,0\\ \end{array}\hskip-3.50006pt\right).$ (68)
@@
$\lim_{x\to 0}\,\,G_{2,4}^{3,1}\left(x\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\,\xi^{2}+1\\[1.49994pt] \xi^{2},\,a,\,b,\,0\\ \end{array}\hskip-3.50006pt\right)\,=\,0.$
@@
For the coherent HD optical channel subjected to GG turbulence ‘without pointing error’ as described in (2), and with perfect CSI at both ends, the exact ASE is given by
@@
$+\frac{1}{\Gamma(a)\Gamma(b)}\,G_{2,4}^{2,2}\left({ab}\mu\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,1\\[1.00006pt] a,b,0,0\\ \end{array}\hskip-3.50006pt\right)\biggr]$ (71)
where $\mu$ is solved from $\int_{\mu}^{\infty}(1/\mu-1/\lambda)f_{{\lambda}}(\lambda)d\lambda\,=\,\mathrm{SNR}$.